OpenStudy (anonymous):

let f(x)=x^2-16. find f^1(x)

OpenStudy (anonymous):

is the equation like this: \[f(x)=x ^{2}-16.\] For the second on is it \[f ^{1}(x) \] or \[f ^{-1}\] or\[f(1)\]?

OpenStudy (mandre):

I think she means the derivative of f(x).

OpenStudy (anonymous):

yes the equation is like that and for the second part its the first one

OpenStudy (mandre):

If it is the derivative then se the product rule on the x^2 part. The derivative of a constant is zero.

OpenStudy (anonymous):

if so, then \[\prime= (2)x ^{2-1}\] 16 is constant.. Derivative of a constant is 0 for x^2 is will be \[n \times ax ^{n-1}\] \[f \prime = 2x\]

OpenStudy (anonymous):

these are the options cuz none of the ones you guys are saying make sence \[4\sqrt{x}\] \[\sqrt{x+16}\] \[\frac{ x^2 }{ 16 }\] \[\frac{ 1 }{ x^2-16}\]

OpenStudy (anonymous):

are you looking for the inverse?

OpenStudy (anonymous):

im guessing so thoses are the options and im confused

OpenStudy (anonymous):

if you're looking for the inverse, then, 1.Switch x and y. We know y=f(x), \[y=x^{2}-16\] \[x=y ^{2}-16\]

OpenStudy (anonymous):

2.solve for y \[y ^{2}= x+16\] \[y=\sqrt{x+16}\] Therefore,\[\sqrt{x+16}\]

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

oh yes i get it i was solving for y wrong

OpenStudy (anonymous):

unfortunately this is an abuse of notation, because in this example \[f^{-1}(x)\] is not a function

OpenStudy (anonymous):

on the second step, you take the square root of both sides, to remove the square in y

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

the very last step should be \[y^2=x+16\] so \[y=\pm\sqrt{x+16}\]

OpenStudy (anonymous):

make sure to include the \(\pm\) in your answer

OpenStudy (anonymous):

@satellite how will I know when to use \[\pm \sqrt{x}\]

OpenStudy (anonymous):

if you have \[a^2=b\] and you are solving for \(a\) it is \(a=\pm\sqrt{b}\)

OpenStudy (anonymous):

for example if \(x^2=9\) then \(x=\pm\sqrt9\) or \(x=3,-3\)

OpenStudy (anonymous):

Does it always have to be like that?

OpenStudy (anonymous):

this is why this is a terrible problem you are asked to find the inverse of a function, but the function is not one to one, so the inverse is not a function written by a lousy math teacher

OpenStudy (anonymous):

but in some books they will consider + as a separate function and - as a separate function

OpenStudy (anonymous):

it would make sense for example if it said "let \(f(x)=x^2-16\) for \(x>4\) then you could find the inverse